Question 3 0.4 pts 3. If M = 9,000, P=6, and Y = 1,500, what is velocity? O a.0.167. O b. 1. O c.4. O d. 36.
Answer #1

B) 1

Here M= 9000,P = 6,Y= 1500

Velocity is the number of times the average dollar is spent to buy goods and services in a given year.

M× V= P× Y

9000× V=6× 1500

9000× V= 9000

V= 9000÷ 9000

V= 1